OpenStudy (anonymous):

What is the mass of 3.98 10-2 moles of zinc chloride (ZnCl2)?

OpenStudy (anonymous):

0ne mole of this salt = 65.3 + 2(35.5) g (check on Zn = 65.3 - im working from memory) then multiply this result by 3.98 * 10^-2

OpenStudy (mertsj):